# A Generalized Theory for Optical Cooling of a Trapped Atom with Spin

### v2

@@ IV Sub-doppler cooling with spin
Doppler cooling alone is often insufficient for achieving ground-state cooling of neutral atoms in optical traps due to two main issues. First, the trapping frequencies in optical traps are typically less than 100 kHz, while the line widths of the D1 and D2 transitions are around 5 MHz, placing them far into the unresolved regime. Second, the ground and excited states have different optical polarizabilities, which can lead to additional heating and atom loss due to far-off-resonant trapping light.
To overcome these limitations, sub-Doppler cooling schemes such as polarization gradient (PG) [13], gray molasses (GM) and $\Lambda$-enhanced GM [17, 18], electromagnetically induced transparency (EIT) cooling [56], and Raman sideband (RSB) [27] employ two-photon transitions between ground-states, as illustrated in Fig. 4(a). By detuning the cooling laser far from the excited states, these techniques mitigate the issues associated with the unresolved regime and differential polarizabilities. Furthermore, the ground hyperfine states in neutral atoms generally experience the same light shifts, resulting in similar trapping frequencies for different spin states.
Two-photon cooling schemes rely on coupling two spin states with different motional quantum numbers, $n$, and energies. The quantum model for PG and GM cooling is depicted in Fig. 4(d,f). In the first stage, a two-photon coherent transition (green) changes the spin and motional states, where the motional state decreases from $n$ to $n-1$. In the second stage, spontaneous emission (red) from the cooling light or a separate beam then re-pumps the atom back to the original spin state, allowing the cooling cycle to repeat. As long as the trapping frequency energy, $\hbar\nu$ is larger than the recoil energy, $E_{\text{recoil}}$, net cooling can be achieved, as the average heating from spontaneous emission is limited to one photon recoil.
In this section, we present a unified spin cooling formalism that encompasses PG, GM, EIT, $\Lambda$-GM, and RSB cooling. We show that all these schemes exhibit similar forms of the effective Hamiltonian and collapse operators in the LD regime and rely on the same principles. PG and GM cooling use a single counter-propagating laser frequency to create light shifts similar to the trapping frequency, as well as the Raman coupling and optical pumping. EIT and $\Lambda$-GM cooling incorporate a second coherent laser to address other spin states and exploit the dark states of a three-level $\Lambda$ system to create darker reservoir states. RSB cooling separates all three processes, using two lasers to couple different ground spin states and a third frequency and beam for optical pumping that uses angular momentum selection rules to keep the cooling state dark.
@@ IV.1 Polarization Gradient Cooling and Gray Molasses
The $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ model is applicable when counter-propagating laser beams with opposite circular polarizations are used. As shown in Fig. 4(b), the resulting polarization is linear at every point in space but rotates along the propagation direction with a period equal to the wavelength. The $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ model relies on the tensor shift and therefore requires a ground-state spin of at least $F=1$. This version is more common because it is also the polarization configuration used by a magneto-optical trap (MOT).
Both forms of PG cooling work by using red-detuned light with a $J\rightarrow J+1$ transition, which is the same configuration used by a MOT for its stretch state cycling condition.
Gray Molasses (GM) is a closely related cooling scheme that operates with blue-detuned light with a $F\rightarrow F$ or $F\rightarrow F-1$ transition. GM has achieved even colder temperatures than PG cooling [58, 17, 59] by storing cold atoms in dark states that do not interact with the cooling light due to polarization and angular momentum selection rules. For example in Fig. 4(e), certain ground-states are not coupled to the excited states. GM works with blue detuned light so that the bright states are higher in energy. Atoms that fall into these dark states are effectively shelved, reducing their interaction with the cooling light. The residual velocity couples the atom back into a bright state, where they can undergo Sisyphus cooling again.
@@ IV.2 Spin cooling theory
First we look at the case of two counter-propagating beams with opposite-handed circular polarized light ($\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ ). The complex polarization is
@@
where $\hat{\bm{e}}_{1,0,-1}$ are the spherical unit vectors $\hat{\bm{e}}_{\pm 1}=\mp\frac{1}{\sqrt{2}}(\hat{\bm{e}}_{x}\pm i\hat{\bm{e}}_{y})$ and $\hat{\bm{e}}_{0}=\hat{\bm{e}}_{z}$, and also The field is Taylor expanded in the last line. A fixed atom sees only linear light.. The intensity is constant everywhere. The polarization is linear but rotates every wavelength, as shown in Fig. 4(b). As it oscillates back and forth in the trap, it sees a small rotation of the field. For convenience, we choose the quantization axis $\hat{\bm{z}}$ to be aligned with linear polarization at equilibrium. The cooling is however independent of the atom position because the light is linear everywhere.
The Hamiltonian and collapse operator in the LD regime are
@@
$+\frac{\Omega}{\sqrt{2}}\left[\hat{D}^{\dagger}_{z}-i\eta\,(\hat{a}+\hat{a}^{\dagger})\hat{D}_{x}^{\dagger}+h.c.\right]$ (61)
@@
The Cartesian dipole operators are defined as $D_{i}=\hat{\bm{i}}\cdot\hat{\bm{D}}$.
The PG cooling processes are illustrated in Fig. 4(d) for a $F=1$ atom. The position-independent interaction $\hat{D}_{z}$ leads to a tensor shift and optical pumping. For the case of $F=1$ ground-state, the population is pumped (red) to the $m=0$ state. The position-dependent term in the interaction $\eta(\hat{a}+\hat{a}^{\dagger})\hat{D}_{x}$ drives an orthogonal polarization transition $m\rightarrow m\pm 1$ while also changing the motional state. This drive is reduced by the LD parameter and produces heating due to spontaneous emission.
The last ingredient for cooling is a two-photon coherent Raman transition (green). One branch of this transition comes from the position-independent term, while the other branch comes from the position-dependent term. As shown in the illustration, together these terms produce a coherent transition from $|m=0,n\rangle$ to $|m=\pm 1,n-1\rangle$. As drawn, the resonance for this two-photon condition is satisfied when the differential light shift between the $m=0$ and $m=\pm 1$ states are equal to the trapping frequency. This process then repeats, cooling the atom.
These processes can be seen more clearly by looking at the effective ground-state operators from Eq. 20. The effective ground-state Hamiltonian is
@@
A spin version of this operator can also be found using Eq. 21.
This spin Hamiltonian is valid for all ground and excited spin combinations. While the ground-state spin determines the spin operators $F_{i}$, the excited state spin only affects the $C^{(i)}$ coefficients. The Hamiltonian and collapse operator both have a position-independent and position-dependent term.
[…]
For circular polarization, the interaction is
$H^{I}_{\text{lin}\perp\text{lin}}(\phi=\pi/2)=\frac{\Omega}{2}\left[\hat{D}_{1}^{\dagger}+\eta(\hat{a}+\hat{a}^{\dagger})\hat{D}_{-1}^{\dagger}+h.c.\right],$ (67)
and for linear polarization,
$H^{I}_{\text{lin}\perp\text{lin}}(\phi=0)=\frac{\Omega}{2}\left[\hat{D}_{z}^{\dagger}+\eta(\hat{a}+\hat{a}^{\dagger})\hat{D}_{x}^{\dagger}+h.c.\right].$ (68)
Before performing the full master equation simulations, our next goal is to develop a simple analytical model for a system with a Hamiltonian and collapse operator in these forms.
Figure 5: PG and GM simulations. (a-b) Clebsch-Gordon coefficients for PG ($F=1$ to $F^{\prime}=1$) and GM ($F=1$ to $F^{\prime}=2$). For PG in linear light, the population pumps to the $m=0$ state, which is also the brightest state. PG cooling requires red-detuned light $(\Delta<0)$. For GM in linear light, the population pumps to $m=0$ which is dark. GM cooling requires blue-detuned light $(\Delta>0)$. (c-d) Simulations of $\langle n\rangle$ versus cooling power for PG and GM. Optimal cooling occurs on a line (light-dashed) where the differential light shift equals the trapping frequency. PG is limited to $\langle n\rangle\approx 1$, while GM cools close to the ground-state due to the dark state. Simulations use $\eta=0.05$ and 8 harmonic states. (e) Optimal cooling for PG and GM versus the Lamb-Dicke parameter $\eta$. Optimal cooling occurs when the differential light shift is equal to the trapping frequency and in the limit of large detuning. The simple model from Eq. 73 is plotted in a lighter color. For PG, $\Delta=-20\Gamma$ and for GM $\Delta=20\Gamma.$ The optical cooling powers were adjusted to make the Raman transition resonant.
We derive an approximate temperature for a simplified model that captures both regimes. We use two spin states $|\!\!\downarrow,n\rangle$ and $|\!\!\uparrow,n\rangle$ and an effective Hamiltonian and collapse operators
@@
$V_{0}$ is a spin-dependent light shift. Position-independent collapse parts with pumping rates $\gamma_{+}$ and $\gamma_{-}$ are the optical pumping between the spin states. The $\Omega_{R}k\hat{x}\hat{F}_{x}$ term produces Raman transitions between $|\uparrow,n\rangle$ and $|\downarrow,n-1\rangle$. $\Omega_{R}=\Omega^{2}/\Delta C^{(2)}$. The optimal light shift for cooling is to make these two states the same frequency by setting $V_{0}=\nu$. The $|\downarrow,n\rangle$ couples to $|\uparrow,n-1\rangle$, but off-resonantly so we can ignore it.
To find an approximate expression for the steady-state temperature, we first use the master equation to determine the steady-state coherences between the Raman-coupled states:
@@
By substituting this coherence into the density matrix equations and ignoring off-resonance coherences, we obtain the transfer rate $\Gamma_{R}$ between $|\uparrow,n\rangle$ to $\downarrow,n-1\rangle$:
@@
$\Delta_{r}$ is the relative detuning between the two states. Optimal cooling occurs when the light shifts bring these states into resonance, i.e., $\Delta_{r}=0$. Although there are off-resonant transitions, such as $|n,\downarrow\rangle\rightarrow|n-1,\uparrow\rangle$, we ignore them in this simple model. By neglecting all off-resonant coherences, we arrive at the population rate equations:
@@
Here $\gamma_{h}=\gamma_{+}+\gamma_{-}$ is a scattering rate related to spontaneous emission heating. Next, we calculate the steady state solution for $\langle n\rangle$. The derivation is as follows. First, we sum the two equations in Eq. 72, yielding a single equation. Next, we multiply the distribution equations by $n$ and sum from $n=0$ to $\infty$, resulting in two equations containing higher-order moments $\langle n^{2}\rangle_{\downarrow}$ and $\langle n^{2}\rangle_{\uparrow}$.
In these four expressions, we have six unknowns: $\langle n^{2}\rangle_{\uparrow}$, $\langle n^{2}\rangle_{\downarrow}$, $\langle n\rangle_{\uparrow}$, $\langle n\rangle_{\downarrow}$, and the fraction of population in either up or down $P_{\uparrow}=\sum_{n=0}P_{\uparrow}(n)$, $P_{\downarrow}=\sum_{n=0}P_{\downarrow}(n)$. We make these equations solvable by simplifying the higher-order moments by assuming that both $P_{\uparrow}(n)$ and $P_{\downarrow}(n)$ follow a geometric distribution.
@@
We can express $\langle n^{2}\rangle_{\uparrow}$ as $\langle n^{2}\rangle_{\uparrow}=2\langle n\rangle_{\uparrow}/P_{\uparrow}+\langle n\rangle_{\uparrow}$, where $P_{\uparrow}$ and $P_{\downarrow}$ appear because they are not a normalized distribution. Then we solve the system of equations for $\langle n\rangle_{\uparrow}+\langle n\rangle_{\downarrow}=\langle n\rangle$.
@@
We then take the limit of large detuning to obtain a simpler analytical expression compared to the complete solution. In the large detuning limit, the cooling rate $\Gamma_{R}=4\Omega_{R}^{2}\eta^{2}/(\gamma_{\scriptscriptstyle+}+\gamma_{\scriptscriptstyle-})$ dominates. The cooling rate $\Gamma_{R}$ is independent of detuning because both $\Omega_{R}^{2}$ and the scattering rate scale as $1/\Delta^{2}$. Consequently, for large detuning, $\Gamma_{R}$ becomes the dominant rate, and the population quickly distributes between the two Raman-coupled states. The asymmetry of the optical pumping then determines the cooling efficiency and the final temperature. In this limit of large detuning, the Raman coupling rate is faster than any scattering rate, and the population simplifies to
$\langle n\rangle=\frac{s}{(1-s)},\quad s=\frac{4\eta^{2}\gamma_{h}+\gamma_{\scriptscriptstyle+}}{\eta^{2}\gamma_{h}+\gamma_{\scriptscriptstyle-}}.$ (73)
We note that a similar expression can also be derived from the no-flow condition solving for $s=P(n+1)/P(n)$ without having to take the expectation values, although other approximations have to be made.
@@
This analytical model is the final result of this model and provides intuition for the cooling performance of PG and GM. It agrees very well with the exact spin 1/2 solution for $\eta<0.1$. For other spin models, it also serves as a fitting function. The ratio $\gamma_{+}/\gamma_{-}$ is then an effective asymmetry of the optical pumping to the lower energy state. This asymmetry is close to unity for PG cooling, and much smaller for cooling with dark state, i.e. GM and $\Lambda$-GM. The ratio $\gamma_{h}/\gamma_{+}$ is an effective scattering rate of the lower energy state. Together with the Lamb-Dicke parameter $\eta$, these three parameters determine the lowest achievable population.
@@
In Fig. 5, we simulate the full master equation with ground and excited states for PG and GM for $F=1$ ground-state in $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ light, meaning that the light is linear at the atom equilibrium position. The PG case is $F=1$ to $F^{\prime}=1$, and the Clebsch-Gordan (CG) coefficients are shown in (a). In the $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ configuration, the $m=0$ state has the largest CG and, consequently, the largest Stark shift. In linear light, however, the population still pumps to the $m=0$ state because of strong diagonal CG coefficients from $|J^{\prime}=2,m=\pm 1\rangle$. Cooling requires that the population is pumped to the lowest energy state. Therefore, cooling requires the red-detuned configuration where the $m=0$, which is the most stark shifted state, is also the lowest energy, as shown in Fig. 4(d).

### v3

@@ IV Sub-doppler cooling with spin
Doppler or sideband cooling alone is often insufficient for achieving ground state cooling of neutral atoms in optical traps due to two main issues. First, the trapping frequencies in optical traps are typically less than 100 kHz, while the line widths of the D1 and D2 transitions are around 5 MHz, placing them far into the unresolved regime. Second, the ground and excited states have different optical polarizabilities, which can lead to additional heating and atom loss.
To overcome these limitations, sub-Doppler cooling schemes such as polarization gradient (PG) [13], gray molasses (GM) and $\Lambda$-enhanced GM [17, 18], electromagnetically induced transparency (EIT) cooling [56], and Raman sideband (RSB) [27] employ two-photon transitions between ground states, as illustrated in Fig. 4(a). By detuning far from the excited states and not populating them, these techniques mitigate the issues associated with the unresolved regime and differential polarizabilities. Furthermore, different ground hyperfine states have the same scalar light shift and no tensor shift, resulting in the same trapping frequencies for different spin states in linear trapping light.
Many of these schemes were developed in the free-space picture, where the atom transverses many wavelengths. In this section, we present a unified spin cooling formalism for trapped atoms that encompasses PG, GM, EIT, $\Lambda$-GM, and RSB cooling. They are all two-photon cooling schemes that rely on coupling two ground spin states with different motional quantum numbers and energies. For example, the free-space and bound pictures for PG and GM are shown in Fig. 4(c-f).
We show that all these schemes exhibit similar forms of the effective Hamiltonian and collapse operators in the LD regime and rely on the same principles. PG and GM cooling use a single laser frequency counter-propagating with a different polarization to create a polarization gradient. The cooling is optimal when the light shifts are similar to the trapping frequency. This same laser is also used to generate the Raman coupling and optical pumping. EIT and $\Lambda$-GM cooling incorporate a second coherent laser resonant to a different ground spin state to create dark states of a three-level $\Lambda$ system. This improves cooling by creating darker reservoir states. RSB cooling separates all three processes, using two lasers and their relative detuning to couple different ground spin states and a third frequency-beam for optical pumping that uses angular momentum selection rules to keep the cooling state dark.
@@ IV.1 Polarization Gradient Cooling and Gray Molasses
[…]
Counter-propagating laser beams with opposite circular polarizations is called the $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ model. As shown in the bottom of Fig. 4(b), the resulting polarization is linear at every point in space but rotates along the propagation direction with a period equal to the wavelength. Because there is no circular light and the resulting vector shift, the $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ model relies on a tensor shift $\Delta E\propto m_{F}^{2}$ and therefore requires a ground state spin of at least $F=1$. The $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ model is much more common experimentally with neutral atoms because it is also the polarization configuration used by a magneto-optical trap (MOT).
[…]
Gray Molasses (GM) is a closely related cooling scheme that operates with blue-detuned light with a $F\rightarrow F$ or $F\rightarrow F-1$ transition. GM has achieved even colder temperatures than PG cooling [58, 17, 59] by storing cold atoms in dark states that do not interact with the cooling light due to polarization and angular momentum selection rules, as shown in Fig. 4(e). Certain ground states are not coupled to the excited states. GM works with blue detuned light so that the bright states are higher in energy. Atoms that fall into these dark states are effectively shelved, reducing their interaction with the cooling light. The residual velocity couples the atom back into a bright state, where they can undergo Sisyphus cooling again.
@@ IV.2 Spin cooling theory
Here we develop the spin cooling master equation, including both the ground and excited states, for PG and GM. First we look at the case of two counter-propagating beams with opposite-handed circular polarized light ($\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ ). For an atom moving along $z$, the complex polarization is
@@
where $\hat{\bm{e}}_{1,0,-1}$ are the spherical unit vectors $\hat{\bm{e}}_{\pm 1}=\mp\frac{1}{\sqrt{2}}(\hat{\bm{e}}_{x}\pm i\hat{\bm{e}}_{y})$ and $\hat{\bm{e}}_{0}=\hat{\bm{e}}_{z}$. The field is Taylor expanded in the last line. We can see from this expression that the atom experiences a rotating linear field, as shown in Fig. 4(b). The intensity is constant everywhere. In the LD regime, As it oscillates back and forth in the trap, it sees a small rotation of the field. For convenience, we now choose a different coordinate system. We choose the quantization axis $\hat{\bm{z}}$ to be aligned with linear polarization at equilibrium. The cooling is however independent of the atom position because the light is linear everywhere.
Substituting the polarization into Eqs. 18 and 19, the Hamiltonian and collapse operators for spin cooling (PG and GM) for $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ in the LD regime are
@@
$+\frac{\Omega}{\sqrt{2}}\left[\hat{D}^{\dagger}_{z}-ik\hat{x}\hat{D}_{x}^{\dagger}+h.c.\right]$ (57)
@@
The Cartesian dipole operators are defined as $\hat{D}_{i}=\hat{\bm{i}}\cdot\hat{\bm{D}}$. For example, the $\hat{D}_{z}$ operator raises all of the ground state spins to the excited state with the same spin $m_{F^{\prime}}=m_{F}$. The operator $\hat{D}_{x}=(\hat{D}_{1}+\hat{D}_{-1})/\sqrt{2}$ creates a superposition of $m_{F^{\prime}}=m_{F}\pm 1$. $\hat{D}_{q}$ for $q=0,\pm 1$ still represents the spherical vector basis.
This master equation both represents PG and GM cooling. For example, Fig. 4(d,f) show the trapped PG and GM processes for an $F=1$ atom. In the trapped situation, the picture for both cases are similar. PG is the case of $F^{\prime}=2$ while is GM is the case of $F^{\prime}=1$. The Clebsch-Gordon for these two cases are also shown in Fig. 6(a) and (c). The position-independent interaction $\hat{D}_{z}$ leads to a tensor shift and optical pumping. In both cases, the population pumps to the $m=0$ state. However, in the PG case the $m=0$ state is bright, whereas in the GM case it is dark. In order to make the $m=0$ state the lower energy, the PG case has to work with red detuning, while the GM case has to work with blue detuning. This is first ingredient for spin cooling: the population is pumped to the lower energy spin state.
The second ingredient for spin cooling is a two-photon Raman transition between the $|m=0,n\rangle$ and $|m=\pm 1,n-1\rangle$, shown in green. This two photon transition arises from the position independent term and the position dependent terms in the interaction. The position-dependent term in the interaction $\hat{x}\hat{D}_{x}$ drives an orthogonal polarization transition $m\rightarrow m\pm 1$ while also changing the motional state. As shown in the illustration, together these terms produce a coherent transition from $|m=0,n\rangle$ to $|m=\pm 1,n-1\rangle$. As drawn, the resonance for this two-photon condition is satisfied when the differential light shift between the $m=0$ and $m=\pm 1$ states are equal to the trapping frequency.
This process then repeats, cooling the atom until the cooling and recoil heating from spontaneous emission reach an equilibrium. As we will see later in the full simulations, GM is much better at preparing ground state population due to the presence of the dark state.
[…]
In this case, however, the position of the atom plays a much more important role because it changes the polarization at equilibrium position For circular polarization, the interaction is
$H^{I}_{\text{lin}\perp\text{lin}}(\phi=\pi/2)=\frac{\Omega}{2}\left[\hat{D}_{1}^{\dagger}+k\hat{x}\hat{D}_{-1}^{\dagger}+h.c.\right],$ (60)
and for linear polarization,
$H^{I}_{\text{lin}\perp\text{lin}}(\phi=0)=\frac{\Omega}{2}\left[\hat{D}_{z}^{\dagger}+k\hat{x}\hat{D}_{x}^{\dagger}+h.c.\right].$ (61)
[…]
Effective spin ground state operators
The cooling and heating processes in the spin model can be seen more clearly by looking at the effective ground state operators with the excited states adiabatically eliminated. Because PG and GM cooling typically use a detuning of ten linewidths or more away from the excited resonance, this is typically a very good approximation. From Eq. 18, the effective ground state Hamiltonian for the $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ case is
@@
A spin version of this operator can also be found using Eq. 22, although it is not written here. These operators are valid for all ground and excited spin combinations. While the ground state spin determines the spin operators $F_{i}$, the excited state spin only affects the $C^{(i)}$ coefficients in the spin representation.
The Hamiltonian and collapse operator both have a position-independent and position-dependent term. As a result, the interpretation becomes similar to the sideband cooling case, where the position independent term drives $n\rightarrow n$, and the position dependent term drives $n\rightarrow n\pm 1$. However, now the operators act within the ground state spin manifold rather than the between the ground and excited states.
Simplified spin model
The ground state spin models accurately predict population transfer, however, they are difficult to use for analytical models. Here we develop a spin 1/2 model, from which we derive simple analytical expressions that we use to understand the final temperature. The spin model is shown in Fig. 5.
Figure 5: Simplified spin model with spins $\uparrow$ and $\downarrow$. The motional states $n$ and $n-1$ are separated by trapping frequency $\nu$. The polarization shifts $|\uparrow,n\rangle$ into resonance with $|\downarrow,n-1\rangle$. The light optical pumps the two states between spins with rates $\gamma_{+}$ and $\gamma_{-}$. The two-photon Raman transition is driven at rate $\Gamma_{R}$. Photon recoil also heats the atom from $n$ to $n+1$ at rate $\gamma_{h}$.
Figure 6: PG and GM simulations. (a-b) Clebsch-Gordon coefficients for PG ($F=1$ to $F^{\prime}=1$) and GM ($F=1$ to $F^{\prime}=2$). For PG in linear light, the population pumps to the $m=0$ state, which is also the brightest state. PG cooling requires red-detuned light $(\Delta<0)$. For GM in linear light, the population pumps to $m=0$ which is dark. GM cooling requires blue-detuned light $(\Delta>0)$. (c-d) Simulations of $\langle n\rangle$ versus cooling power for PG and GM. Optimal cooling occurs on a line (light-dashed) where the differential light shift equals the trapping frequency. PG is limited to $\langle n\rangle\approx 1$, while GM cools close to the ground state due to the dark state. Simulations use $\eta=0.05$ and 8 harmonic states. (e) Optimal cooling for PG and GM versus the Lamb-Dicke parameter $\eta$. Optimal cooling occurs when the differential light shift is equal to the trapping frequency and in the limit of large detuning. The simple model from Eq. 72 is plotted in a lighter color. For PG, $\Delta=-20\Gamma$ and for GM $\Delta=20\Gamma.$ The optical cooling powers were adjusted to make the Raman transition resonant.
We assume two spin states $|\!\!\downarrow,n\rangle$ and $|\!\!\uparrow,n\rangle$. We use an effective Hamiltonian and collapse operators that are similar in form to Eqs. 63 and 64. We represent the light shift with $V_{0}$ and Raman coupling with $\Omega_{R}$ in the Hamiltonian and collapse operators
@@
[…]
Optimal cooling occurs when the light shift is equal to the trapping frequency $\nu$, so that $|\uparrow,n\rangle$ and $|\downarrow,n-1\rangle$ are the same energy. Then the Raman transition is resonant. The collapse operators also contain a position dependent term that couples different motional states and represents recoil heating.
Now we derive an approximate expression for the steady-state temperature of this model. First we estimate the cooling rate due to the Raman transition. The position dependent Hamiltonian coherently couples $|\uparrow,n\rangle$ to $|\downarrow,n-1\rangle$. However, this alone would not result in cooling because the population would Rabi oscillate back and forth. The optical pumping rates $\gamma_{+}$ and $\gamma_{-}$ produce decoherence that results in population transfer. We can estimate this transfer rate by first solving for the steady state coherence between the Raman-coupled states using the master equation for the density matrix and setting the time derivative to zero:
@@
By substituting this coherence into the density matrix equations and ignoring off-resonance coherences, we obtain the transfer rate $\Gamma_{R}$ between $|\uparrow,n\rangle$ to |||| $\downarrow,n-1\rangle$:
@@
$\Delta_{r}$ is the relative detuning between the two Raman coupled states. Optimal cooling occurs when the light shift matches the trapping frequency and brings these states into resonance, i.e., $\Delta_{r}=0$. Although there are off-resonant transitions, such as $|\downarrow,n\rangle\rightarrow|\uparrow,n-1\rangle$, we ignore them. By neglecting all off-resonant coherences, we arrive at the population rate equations:
@@
We have defined a total spontaneous emission rate
$\gamma_{h}=\gamma_{+}+\gamma_{-}.$
Next, we calculate the steady state solution for $\langle n\rangle$. The derivation is as follows. First, we perform the sum $\sum_{n=0}^{\infty}$ to the two equations in Eq. 68. Next we multiply these two equations by $n$ and perform the same sum, yielding two more equations. This results in a total of four equations. We set the time derivatives to zero to calculate the steady-state. We then have six unknowns in these four equations: $\langle n^{2}\rangle_{\uparrow}$, $\langle n^{2}\rangle_{\downarrow}$, $\langle n\rangle_{\uparrow}$, $\langle n\rangle_{\downarrow}$, and the fraction of population in either up or down
$P_{\uparrow}=\sum_{n=0}P_{\uparrow}(n),\quad P_{\downarrow}=\sum_{n=0}P_{\downarrow}(n).$ (69)
Two of these terms are higher-order moments. If we assume that $P_{\uparrow}(n)$ and $P_{\downarrow}(n)$ follow a geometric distribution, then we can use properties of the geometric series to write them as
$\langle n^{2}\rangle_{\uparrow}=2\langle n\rangle_{\uparrow}/P_{\uparrow}+\langle n\rangle_{\uparrow}$ (70)
$\langle n^{2}\rangle_{\downarrow}=2\langle n\rangle_{\downarrow}/P_{\downarrow}+\langle n\rangle_{\downarrow}.$ (71)
@@
Here $P_{\uparrow}$ and $P_{\downarrow}$ appear because $P_{\uparrow}(n)$ and $P_{\downarrow}(n)$ are not a normalized distribution as they do not sum up to one.
@@
We then have four equations and four unknowns. Then we solve the system of equations for $\langle n\rangle_{\uparrow}+\langle n\rangle_{\downarrow}=\langle n\rangle$. This expression is long, but we can find a simpler expression by taking the limit of large detuning $\Delta$. There are two main processes in the time evolution. The Raman transfer $\Gamma_{R}$ and the pumping rates $\gamma_{\pm}$. In the large detuning limit, the cooling rate $\Gamma_{R}=4\Omega_{R}^{2}\eta^{2}/(\gamma_{\scriptscriptstyle+}+\gamma_{\scriptscriptstyle-})$ dominates. The pumping rates scale as $1/\Delta^{2}$. The cooling rate $\Gamma_{R}$ is independent of detuning because both $\Omega_{R}^{2}$ and the scattering rate scale as $1/\Delta^{2}$. Consequently, for large detuning, $\Gamma_{R}$ becomes the dominant rate, and the population quickly distributes between the two Raman-coupled states. The asymmetry of the optical pumping then determines the cooling efficiency and the final temperature.
In this limit of large detuning, the Raman coupling rate is faster than any scattering rate, and the steady-state population simplifies to
$\langle n\rangle=\frac{s}{1-s},\quad s=\frac{4\eta^{2}\gamma_{h}+\gamma_{\scriptscriptstyle-}}{\eta^{2}\gamma_{h}+\gamma_{\scriptscriptstyle+}}.$ (72)
This expression for the steady-state population is the final result of this model. It agrees very well with the exact spin 1/2 solution for $\eta<0.1$. We can look at a few limiting cases. The ratio $\gamma_{-}/\gamma_{+}$ is the pumping asymmetry between the spin states. For PG cooling this ratio is approximately one, as there is a small asymmetry. The factor $\eta^{2}\gamma_{h}$ is a spontaneous emission rate that leads to heating. For PG and small $\eta$, the steady-state population becomes $s\approx\gamma_{-}/\gamma_{+}$, which gives $\langle n\rangle\approx 1$. And this does not depend on the LD parameter.
Experimentally, both PG and GM operate in a low B-field regime. Since they use light shifts to adjust the motional trap frequencies into resonance, the Zeeman shifts must remain smaller than the trapping frequency, providing an estimate for how small the B-field needs to be.
For the GM case though, the lower energy state is dark. For $\gamma_{-}=0$, then $\langle n\rangle\approx 4\eta^{2}\frac{\gamma_{h}}{\gamma_{+}}$. Because $\gamma_{h}=\gamma_{+}+\gamma_{-}$, this ratio is close to unity. This scales with with the LD parameter squared and can be made much smaller than in the PG case.
In the derivation of the steady-state temperature for the simplified model, we ignore the off-resonant coherent coupling between $|\uparrow,n-1\rangle$ and $|\downarrow,n\rangle$. From Eq. 67 and using the off-resonant detuning $\Delta_{r}=2\nu$, we obtain a transfer rate between these states of
$\Gamma_{R}^{\text{off}}\approx\frac{\eta^{2}\Omega_{R}^{2}\gamma_{h}}{4\nu^{2}}.$
The resonance condition is close to $\Omega_{R}\approx\nu$, so we can eliminate those terms and get an approximate heating rate from off-resonant scattering:
$\Gamma_{R}^{\text{off}}\approx\frac{\eta^{2}\gamma_{h}}{4}.$
This heating rate is four times less than the heating rate $\eta^{2}\gamma_{h}$, which arises from the position-dependent collapse operator in this model, and can thus be ignored. In the effective model, we can consider this off-resonant coherent transfer as being incorporated into the effective $\gamma_{h}$.
@@
In the next section we perform full simulations and compare it to the model. For other $F$, this model will also serve as a useful fitting function. The ratio $\gamma_{-}/\gamma_{+}$ is then an effective asymmetry of the optical pumping to the lower energy state. This asymmetry is close to unity for PG cooling, and much smaller for cooling with dark state, i.e. GM and $\Lambda$-GM. The ratio $\gamma_{h}/\gamma_{+}$ is an effective scattering rate of the lower energy state. Together with the Lamb-Dicke parameter $\eta$, these three parameters determine the lowest achievable population.
Simulations
@@
In Fig. 6, we simulate the full master equation with ground and excited states for PG and GM for $F=1$ ground state in $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ light, meaning that the light is linear at the atom equilibrium position. The PG case is $F=1$ to $F^{\prime}=1$, and the Clebsch-Gordan (CG) coefficients are shown in (a). In the $\sigma_{\scriptscriptstyle+}\!-\!\sigma_{\scriptscriptstyle-}$ configuration, the $m=0$ state has the largest CG and, consequently, the largest Stark shift. In linear light, however, the population still pumps to the $m=0$ state because of strong diagonal CG coefficients from $|J^{\prime}=2,m=\pm 1\rangle$. Cooling requires that the population is pumped to the lowest energy state. Therefore, cooling requires the red-detuned configuration where the $m=0$, which is the most stark shifted state, is also the lowest energy, as shown in Fig. 4(d).